OpenStudy (anonymous):

The function f(t)=5cos(pi/4^t)+11 represents the tide in the Dark Sea. It has a maximum of 16 feet when time (t) is 0 and a minimum of 6 feet. The sea repeats this cycle every 8 hours. After 6 hours, how high is the tide?

OpenStudy (anonymous):

A. 11 feet B. 6 feet C. 8.5 feet D. 10.5 feet

OpenStudy (perl):

you can plug in 6 into t

OpenStudy (anonymous):

\[f(6)=5\cos (\pi/4^6)+11 l\]

OpenStudy (anonymous):

like that?

OpenStudy (perl):

thats a bit strange to have time as the exponent

OpenStudy (anonymous):

Yeah I'm not really how to solve it now. The word problems on my homework are always hard.

OpenStudy (perl):

$$ \Large{ f(t)=5\cos(\frac{\pi}{4}\cdot t)+11 \\ ~\\ f(6) = 5\cos(\frac{\pi}{4}\cdot 6)+11 }$$

OpenStudy (anonymous):

How do I go about solving this?

OpenStudy (perl):

we can use a calculator or our knowledge of the unit circle

OpenStudy (perl):

what is cos (6/4 pi ) ?

OpenStudy (perl):

simplify to cos (3/2 pi )

OpenStudy (anonymous):

Then what?

OpenStudy (perl):

multiply that by 5 , then add 11

OpenStudy (perl):

we can use google's calculator https://www.google.com/search?q=5cos%286%2F4+*pi%29+%2B11&ie=utf-8&oe=utf-8

OpenStudy (anonymous):

I got 11 feet

OpenStudy (perl):

correct ☺

OpenStudy (anonymous):

Could you help with 2 other questions please?

OpenStudy (perl):

ok

OpenStudy (anonymous):

Simplify (sinƟ-cosƟ)^2+(sinƟ+cosƟ)^2 A. -sin^2Ɵ B. -cos^2Ɵ C. 0 D. 2

OpenStudy (perl):

i would expand that

OpenStudy (perl):

(sinƟ-cosƟ)^2+(sinƟ+cosƟ)^2 =(sinΘ-cosΘ)(sin Θ - cos Θ ) + (sinΘ+cosΘ)( sin Θ + cos Θ)

OpenStudy (anonymous):

Don't they all cross out?

OpenStudy (perl):

not necessarily, did you foil it out?

OpenStudy (anonymous):

One sec

OpenStudy (anonymous):

I'm confused on how to use foil on it

OpenStudy (perl):

(sinƟ-cosƟ)^2+(sinƟ+cosƟ)^2 =(sinΘ-cosΘ)(sin Θ - cos Θ ) + (sinΘ+cosΘ)( sin Θ + cos Θ) = sin^2 Θ - cosΘ sin Θ - cosΘsin Θ + cos^2 Θ + sin^2 Θ + cosΘ sin Θ + cosΘ sin Θ + cos^2Θ

OpenStudy (anonymous):

I got \[\sin ^2\theta-\cos^2\theta+\sin^2\theta+\cos^2\theta\]

OpenStudy (perl):

(sinƟ-cosƟ)^2+(sinƟ+cosƟ)^2 =(sinΘ-cosΘ)(sin Θ - cos Θ ) + (sinΘ+cosΘ)( sin Θ + cos Θ) = sin^2 Θ - cosΘ sin Θ - cosΘsin Θ + cos^2 Θ + sin^2 Θ + cosΘ sin Θ + cosΘ sin Θ + cos^2Θ =sin^2 Θ - 2cosΘ sin Θ+ cos^2 Θ + sin^2 Θ +2cosΘ sin Θ + cos^2Θ =2sin^2 Θ + 2cos^2 Θ = 2 ( sin^2 Θ + cos^2 Θ ) = 2*1

OpenStudy (anonymous):

Oh I understand it now. It's not that hard after it's explained lol

OpenStudy (perl):

☺ ☺

OpenStudy (anonymous):

So it's 2. I have one more question that I didn't understand

OpenStudy (anonymous):

Compare the following functions. Which function has the smallest minimum?

OpenStudy (anonymous):

Thats one of them hang on

OpenStudy (anonymous):

f(x)= -5sin(2x-pi)+2 h(x) x y -2 10 -1 7 0 5 1 3 2 5 3 7 4 10

OpenStudy (anonymous):

The graph is g(x)

OpenStudy (anonymous):

@perl still there?

OpenStudy (anonymous):

I think it is f(x)

OpenStudy (perl):

what is the minimum of f(x)

OpenStudy (anonymous):

I got -3

OpenStudy (perl):

It looks like -2 to me

OpenStudy (anonymous):

So they all have the same minimum?

OpenStudy (perl):

wait, there are three functions?

OpenStudy (anonymous):

Yes, the graph, the f(x) function and the table

OpenStudy (perl):

ok the minimum of the function in the table is 3. the minimum of the graph is -2 the minimum of the sine function is -5*1 + 2 = -3

OpenStudy (anonymous):

So it is f(x)

OpenStudy (perl):

correct :)

OpenStudy (anonymous):

Yay! thank you so much! :)

OpenStudy (anonymous):

I fanned :)

OpenStudy (perl):

Thanks (• ◡•)|